Vernon Fruit Drink Company planned to make 209,000 containers of apple juice. It expected to use two cups of frozen apple concentrate to make each container of juice, thus using 418,000 cups of frozen concentrate. The standard price of one cup of apple concentrate is $0.25. Vernon actually paid $123,627 to purchase 426,300 cups of concentrate, which was used to make 210,000 containers of apple juice.

Required:

  1. Compute the actual price per cup of concentrate. (Round your answer to 2 decimal places.)

  2. Compute the standard quantity (number of cups of concentrate) required to produce the containers.

  3. Compute the materials price variance and indicate whether it is favorable (F) or unfavorable (U). (Select "None" if there is no effect (i.e., zero variance). Round "Price variance" to 2 decimal places.)

  4. Compute the materials usage variance and indicate whether it is favorable (F) or unfavorable (U). (Select "None" if there is no effect (i.e., zero variance). Round "Usage variance" to 2 decimal places.)
Answer #1

b) Actual price per cup = 123627/426300 = 0.29

c) Standard quantity = 210000*2 = 420000

d) Material price variance = (0.25-0.29)*426300 = 17052 U

e) Material usage variance = (420000-418000)*0.25 = 500 F